# Role of defects in atom probe analysis of sol-gel silica

### v1

@@ Abstract
Silica is a suitable material to encapsulate proteins at room temperature, enabling their analysis at the atomic level using laser-assisted atom probe tomography (La-APT). In this study, we show that UV and deep-UV lasers can achieve a high success rate in La-APT of silica in terms of chemical resolution and three-dimensional image volume, with both lasers providing comparable results. Since the La-APT analyses are driven by photon absorption, in order to understand the mechanisms behind the enhanced absorption of UV light, we performed density functional theory calculations to model the electronic and optical properties of amorphous silica matrices generated using a Monte Carlo approach to structural optimisation. In particular, we have investigated the role of various defects introduced during sample preparation, such as substitutional and interstitial carbon, sodium and gallium ions, and hydrogen. Our results show that the presence of defects increases the absorption of silica in the UV and deep-UV range and thus improves the La-APT capabilities of the material. However, due to the low density of free charge carriers resulting from the absorption of laser energy by defects, deviations from the nominal chemical composition and suboptimal chemical resolution may occur, potentially limiting the optimal acquisition of APT mass spectra.
@@ 1 Introduction
Silicon dioxide (SiO2), commonly known as silica, is the most abundant element in the earth’s crust 1, 2. It occurs in nature in various forms, such as sand, minerals and quartz. Silica has been processed and used by humans for a long time, especially in the production of glass. At the molecular level, silica is structured as an interconnected network of silicon atoms that are covalently bonded in a tetrahedral coordination by oxygen bridges shared by two silicon tetrahedra 3, 4. In amorphous silicon dioxide, this local, short-range order does not continue into a longer-range order, but forms an amorphous material with characteristic properties such as transparency to visible light, brittleness and chemical inertness.
@@
Silica has been synthesised for various purposes using different methods. One commonly used method is the Stöber process, in which a silicate precursor is hydrolysed under either alkaline (pH $\simeq$ 11-12) or basic conditions using ammonia as a catalyst and then condensed to form the silica network 5. Another method is the sol-gel process, in which a precursor solution, such as sodium silicate with a high pH value, forms a macroscopic gel by polymerisation through condensation of the silicate monomers when the pH value is lowered. Sol-gel silica is particularly useful for encapsulating proteins in a solid material that retains their native structure at room temperature similar to an aqueous environment 6. This allows proteins to be analysed at the atomic level 7. For example, Immunoglobulin G (IgG) proteins embedded in silica were imaged three-dimensionally near atomic resolution using laser-assisted atom probe tomography (La-APT) and provided with information about their elemental composition 8.
@@
Atom probe tomography (APT) is a powerful technique for material characterisation based on the controlled field evaporation of individual ions from a needle-shaped specimen. These ions move along the applied field and are detected by a 2D position-sensitive detector 9, 10. In La-APT, the process of field ion emission is triggered by ultrafast laser pulses 11. In particular, a combination of femtosecond laser pulses and static fields is used to briefly heat the sample to reduce the energy barrier for field ion evaporation 12, 13, 14. Time-of-flight mass spectrometry enables the chemical identification of the ions, while the position of the impact on the detector is used to reconstruct the initial position and thus generate the 3D image of the sample by back-projection 9, 10.
@@
While silica has already been used as a framework for the encapsulation of biomolecular substrates, where the native biomolecule can be preserved by minimising mechanical stresses, it poses some major challenges for APT analysis. Indeed, the poor conductivity of silica represents a problem for APT analysis as the applied potential cannot be easily carried to the apex of the sample to obtain the electrostatic field necessary for field evaporation 15, 16, 17. In this context, recent experimental and theoretical studies have shown that THz pulses with negative polarity can trigger the evaporation of cations from nanoneedle-shaped amorphous silica specimens efficiently 14. In addition, in a previous work 8 a green laser with a wavelength of 532 nm (corresponding to a photon energy of 2.4 eV) was applied to a silica sample, resulting in a low evaporation yield and a small analysed volume due to the fracture of the sample.
@@
The main reason for the difficulties in La-APT analysis of silica using a green laser is related to its low absorption of green light, since the band gap of silica is about 9 eV, which is higher than the photon energy of the green laser 18. However, defects are frequently observed in sol-gel silica and can also occur during sample preparation of the nano-needle 19, 20. Defects can increase the sub-band gap light absorption and thus the ion yield of La-APT analysis.
@@
In this work, we have performed an experimental investigation of silica samples synthesised with our sol-gel method using La-APT devices equipped with green, ultraviolet (UV) and deep-UV lasers at 515, 343 and 258 nm, combined with a computational study of the electronic and optical absorption properties of silica based on density functional theory (DFT) simulations that account for the presence of defects within the silica matrix. DFT is widely regarded as the standard computational method to study the electronic structure of solids, although more sophisticated approaches such as the GW approximation 21, 22, 23, 24 or the Bethe-Salpeter equation 25 provide improved quasiparticle descriptions and thus more accurate optical absorption spectra. However, their application to the class of materials investigated in this study, which are disordered and amorphous, poses significant methodological and computational challenges, making DFT a favourable compromise between accuracy and computational cost. We show that the defects in silica most likely increase the absorption of UV and deep-UV light. This explains why we report successful La-APT analyses with high yield and large analysed volume when we use light in this frequency range, but much less often when we use green light. Nevertheless, the experimental results also show that the density of charge carriers generated by the absorption of laser energy is not sufficient to fully correct the problem of low conductivity in La-APT analyses of insulating materials in the UV and deep-UV range.
@@ 2 Materials and Methods
2.1 Samples preparation
The silica samples were prepared using a sodium silicate solution (Sigma-Aldrich) as starting material, as shown in Figure 1a). The sodium silicate solution was diluted 1:3 in water and gelation was initiated by passing the solution through a syringe filled with a Dowex 50WX ion exchange resin (Sigma-Aldrich) activated with HCl [1M] to adjust the pH from alkaline to neutral, according to the procedure described in Refs. 8, 26. After activation with HCl, the ion exchange resin was adjusted by passing Milli-Q water through it in small increments until an extruded pH value of around 5 was reached. The diluted sodium silicate solution was then passed through the syringe to obtain a neutral pH. A solid glass was formed by drying the silica gel at 37 °C overnight.
@@ 2.1 Samples preparation
The specific surface area and pore size distribution of the synthesised silica samples were determined by nitrogen sorption measurements using an ASAP 2020 instrument (Micromeritics Instrument Corporation, Norcross GA). The specific surface area was assessed using the Brunauer-Emmett-Teller (BET) method 27, the micropores using the Horvath-Kawazoe method 28 and the mesopores using the Barrett-Joyner-Halenda (BJH) sorption method 29. The measurements resulted in a specific surface area of 587 m2/g for a representative silica sample. The material contains both micropores and mesopores. The micropores have an average width of 6.8 Å and the mesopores have an average size of 40 Å. Further details on the isothermal measurements can be found in the Supplementary Information (SI, see Figure S.1).
@@
Finally, the APT specimens were prepared with the conventional focused ion beam scanning electron microscopy (FIB-SEM) in-situ lift-out protocol using a FIB-SEM dual beam system (FEI Versa 3D) with a single isotope Ga+ source (69 amu) 30, 31. An area of interest on the sample surface was identified and covered with a 2$\times$20 $\mu$m Pt protective strip deposited with the device’s gas injection system (GIS) using a (methyl-cyclopentadienyl)-trimethyl platinum precursor gas. A cantilever was prepared for lift out by milling trenches with the stage tilted 30° with respect to the ion beam. The cantilever was lifted out with an Omniprobe micromanipulator and placed in segments on a flat silicon or tungsten rod. The lift-out segments were polished with the focussed ion beam (FIB) by annular milling to obtain the needle-shaped sample shown in Figure 1b,c. We notice that the high charging effect due to the low electrical conductivity of silica reduces the image resolution.
@@ 2.2 Atom Probe Tomography
The LEAP 6000 XR operated in laser mode (wavelength 257.5 nm, $\approx$4.8 eV, deep-UV) with laser pulse energies (LPE) between 100 and 250 pJ focussed on a spot of 2 $\mu$m diameter, which corresponds to an energy density between 30 and 80 J/m2. The sample temperature was set to either 35 K or 50 K. The pulse rate was set using the device’s automatic pulse rate control so that ions could be detected up to 250 Da before the next pulse, regardless of the voltage. This meant that the pulse rate increased with increasing voltage due to shorter flight times.
@@ 2.3 Theoretical and computational methods
The amorphous silica matrices are modelled by assuming an ideal $\beta$-cristobalite structure with a density of $0.0662$ atoms/Å3. The Monte Carlo approach developed by Wooten, Winer and Weaire 32 is used to generate defect-free amorphous structures of SiO2. In this method, the structure of the material is represented as a network of interconnected atoms and its equilibrium configuration is found by Monte Carlo moves acting on such initial structures by displacing the atoms 33. The acceptance (rejection) criterion of the atomic displacement is chosen according to the Boltzmann probability distribution, where the decrease (increase) in potential energy is evaluated using a force field parameterised for SiO2 34. The resulting defect-free minimum-energy amorphous silica matrices are contained in a cubic box with periodic boundary conditions. In particular, we have generated three different sizes of simulation cells corresponding to a cubic box with an edge of $L=10.283$, $L=14.26$ and $L=21.39$ (see Figure S.4 of the SI for an image of one of the structures obtained).
The optical properties of these systems are evaluated in the framework of linear response theory by calculating the imaginary part of the complex dielectric tensor $\epsilon^{(2)}_{\alpha,\beta}$ (where $\alpha,\beta$ denotes the tensor components with respect to a set of Cartesian axes), whose spatial average is linearly related to the absorption coefficient (modulo some weighting factors, see Figures S.7 and S.8 of the SI for a plot of the real and imaginary parts of $\epsilon(\omega)$ and the refractive index, which is in the range 1.5-1.7 in the energy region of interest). In our calculations, local field effects (LFE) are neglected and the expression for $\epsilon^{(2)}_{\alpha,\beta}$ is given by 23, 35
@@
where $e$ is the electron charge, $\Omega$ is the volume of the box, $N_{\mathbf{k}}$ is the number of points used to describe the electronic bands, $m$ is the mass of the electron, $\hat{\mathbf{M}}$ is the squared matrix element that connects via the radiation perturbative field the initial and final wave functions of the electrons, which belong to different bands labelled $n$ and $n^{\prime}$ but have the same crystal momentum $\mathbf{k}$, while $E_{n,\mathbf{k}}$ is the energy value of band $n$ at the point $\mathbf{k}$ of reciprocal space. Eq. 1 is obtained under the assumption that the transmitted momentum can be neglected (long wavelength limit), but can be extended to finite momentum transfer. The $\delta$ functions are slightly smeared to avoid unphysical excitations with infinite lifetimes. When comparing the spectra of the three boxes with different edges, we found that convergence is achieved with the box with $L=14.26$, which corresponds to 192 atoms (see Figure S.5 in the SI, where we compare the results for $L=14.26$ and $L=21.42$).
@@
The calculation of $\epsilon^{(2)}_{\alpha,\beta}$ was performed using DFT as implemented in the Quantum Espresso code suite 36. The ground state electron density was determined with a plane wave cutoff equal to $80$ Rydberg with the Perdew-Burke-Ernzerhof (PBE) functional. Using the same functional, we also investigated the effect of a finite constant electric field on the electronic properties of the pristine matrices and found negligible effects for all amorphous silica configurations (see Figure 7b). Once PBE self-consistency is achieved, we improved the estimated electronic bandgap by using the HSE (Heyd–Scuseria–Ernzerhof) hybrid exchange–correlation functional 37 (see Figure S.6 in SI), which indeed yields a value of about 8 eV, which is much closer to the experimental values in the range of 8 to 9 eV 38, 39, 40.
@@
Recall that both sol-gel synthesis and specimen preparation for APT analysis usually result in defects in the pristine amorphous silica matrix, such as vacancies, dislocations and interstitials. In particular, the use of FIB for lift-out can lead to substitution or displacement of atomic constituents by the implantation of Ga+ ions, which can lose all their kinetic energy through elastic and inelastic interactions in the sample and be adsorbed. In the case of gallium, depending on the ion concentration, this can lead to significant changes in the structural and optical properties of the sample. To determine the effects of such defects on the optical properties of the sample, we performed the calculation of the imaginary part of the dielectric function in the long wavelength limit (see equation 1) for a number of typical defect configurations by including interstitial hydrogen, carbon and sodium atoms, oxygen vacancies 21, interstitial and substitutional gallium ions at different concentrations. The defective silica structures were optimised and the optical properties for each configuration were evaluated using DFT with the HSE exchange-correlation functional, which ensures a more accurate evaluation of the resulting absorption properties.
3 Results and Discussion

### v2

@@ Abstract
Silicon dioxide is a suitable material to encapsulate proteins at room temperature so that they can be analysed at the atomic level using laser-assisted atom probe tomography (La-APT). To achieve this goal, in this study we show that UV and deep UV lasers can achieve a high success rate in La-APT of silica in terms of chemical resolution and three-dimensional image volume, with both lasers providing comparable results. Since the La-APT analyses are driven by photon absorption, in order to understand the mechanisms behind the enhanced absorption of UV light, we performed density functional theory calculations to model the electronic and optical properties of amorphous silica matrices generated using a Monte Carlo approach to structural optimisation. In particular, we have investigated the role of various defects introduced during sample preparation, such as substitutional and interstitial carbon, sodium and gallium ions, and hydrogen. Our results show that the presence of defects increases the absorption of silica in the UV and deep-UV range and thus improves the La-APT capabilities of the material. However, due to the low density of free charge carriers resulting from the absorption of laser energy by defects, deviations from the nominal chemical composition and suboptimal chemical resolution may occur, potentially limiting the optimal acquisition of APT mass spectra.
@@ 1 Introduction
Silicon dioxide (SiO2), commonly known as silica, is the most abundant element in the earth’s crust 1, 2. It occurs in nature in various forms, e.g. as sand, minerals and quartz. At the molecular level, most silica allotropes are structured as an interconnected network of silicon atoms covalently bonded in tetrahedral coordination by oxygen bridges shared by two silicon tetrahedra 3, 4 (however, not all allotropes are tetrahedral, e.g. stishovite, which forms at very high pressure, is a polymorph in which each Si is surrounded by 6 oxygen atoms).
In amorphous silicon dioxide, this local, short-range order does not continue into a longer-range order, but forms an amorphous material with characteristic properties such as transparency to visible light, brittleness and chemical inertness.
@@
Silica has a long history of being processed and utilized, especially in glass manufacturing, and has also been synthesised using different methods. One commonly used method is the Stöber process, in which a silicate precursor is hydrolysed under either alkaline (pH $\simeq$ 11-12) or basic conditions using ammonia as a catalyst and then condensed to form the silica network 5. Another method is the sol-gel process, in which a precursor solution, such as sodium silicate with a high pH value, forms a macroscopic gel by polymerisation through condensation of the silicate monomers when the pH value is lowered. Sol-gel silica is particularly useful for encapsulating proteins in a solid material that retains their native structure at room temperature similar to an aqueous environment 6, 7. This allows proteins to be analysed at the atomic level 8. For example, Immunoglobulin G (IgG) proteins embedded in silica were imaged three-dimensionally with near-atomic resolution using laser-assisted atom probe tomography (La-APT), which provided information about their elemental composition 9.
@@
Atom probe tomography (APT) is a powerful technique for material characterisation based on the controlled field evaporation of individual ions from a needle-shaped specimen. These ions move along the applied field and are detected by a 2D position-sensitive detector 10, 11. In La-APT, the process of field ion emission is triggered by ultrafast laser pulses 12. In particular, a combination of femtosecond laser pulses and static fields is used to briefly heat the sample to reduce the energy barrier for field ion evaporation 13, 14, 15. Time-of-flight mass spectrometry enables the chemical identification of the ions, while the position of the impact on the detector is used to reconstruct the initial position and thus generate the 3D image of the sample by back-projection 10, 11.
@@
While silica has already been used as a framework for the encapsulation of biomolecular substrates, where the native biomolecule can be preserved by minimising mechanical stresses, it poses some major challenges for APT analysis. Indeed, the very low conductivity of amorphous silica, which is an electrical insulator but whose conductivity can increase slightly under certain conditions (impurities, doping, high temperatures or radiation) represents a challenge for APT analysis, as the applied potential cannot easily be carried to the apex of the sample to obtain the electrostatic field required for field evaporation. 16, 17, 18. In this context, recent experimental and theoretical studies have shown that THz pulses with negative polarity can trigger the evaporation of cations from nanoneedle-shaped amorphous silica specimens efficiently 15. In addition, in a previous work 9 a green laser with a wavelength of 532 nm (corresponding to a photon energy of 2.4 eV) was applied to a silica sample, resulting in a low evaporation yield and a small analysed volume due to the fracture of the sample.
@@
The main reason for the difficulties in La-APT analysis of silica using a green laser is related to its low absorption of green light, since the band gap of silica is about 9 eV, which is higher than the photon energy of the green laser 19. However, defects are frequently observed in sol-gel silica and can also occur during sample preparation of the nano-needle 20, 21. Defects can increase the sub-band gap light absorption and thus the ion yield of La-APT analysis.
@@
In this work, we have performed an experimental investigation of silica samples synthesised with our sol-gel method using La-APT devices equipped with green, ultraviolet (UV) and deep-UV lasers at 515, 343 and 258 nm, combined with a computational study of the electronic and optical absorption properties of silica based on density functional theory (DFT) simulations that account for the presence of defects within the silica matrix. DFT is widely regarded as the standard computational method to study the electronic structure of solids, although more sophisticated approaches such as the GW approximation 22, 23, 24, 25, 26, 27 or the Bethe-Salpeter equation 28, 29 provide improved quasiparticle descriptions and thus more accurate optical absorption spectra. However, their application to the class of materials investigated in this study, which are disordered and amorphous, poses significant methodological and computational challenges, making DFT a favourable compromise between accuracy and computational cost. We show that the defects in silica most likely increase the absorption of UV and deep-UV light. This explains why we report successful La-APT analyses with high yield and large analysed volume when we use light in this frequency range, but much less often when we use green light. Nevertheless, the experimental results also show that the density of charge carriers generated by the absorption of laser energy is not sufficient to fully correct the problem of low conductivity in La-APT analyses of insulating materials in the UV and deep-UV range. The aim of this work is in particular to determine the specific role of defects in the light absorption of silica in order to isolate their influence from other effects in APT experiments, such as that due to heating.
@@ 2 Materials and Methods
2.1 Sample preparation
The silica samples were prepared using a sodium silicate solution (Sigma-Aldrich) as starting material, as shown in Figure 1a. The sodium silicate solution was diluted 1:3 in water and gelation was initiated by passing the solution through a syringe filled with a Dowex 50WX ion exchange resin (Sigma-Aldrich) activated with HCl [1M] to adjust the pH from alkaline to neutral, according to the procedure described in Refs. 9, 30. After activation with HCl, the ion exchange resin was adjusted by passing Milli-Q water through it in small increments until an extruded pH value of around 5 was reached. The diluted sodium silicate solution was then passed through the syringe to obtain a neutral pH. A solid glass was formed by drying the silica gel at 37 °C overnight.
@@ 2.1 Sample preparation
The specific surface area and pore size distribution of the synthesised silica samples were determined by nitrogen sorption measurements using an ASAP 2020 instrument (Micromeritics Instrument Corporation, Norcross GA). The material actually contains both micropores and mesopores. The specific surface area was assessed using the Brunauer-Emmett-Teller (BET) method 31, the micropores with the Horvath-Kawazoe method 32 and the mesopores with the Barrett-Joyner-Halenda (BJH) sorption method 33. The measurements resulted in a specific surface area of 587 m2/g for a representative silica sample. The micropores have an average width of 6.8 Å and the mesopores have an average size of 40 Å. Further details on the isothermal measurements can be found in the Supplementary Information (SI, see Figure S.1).
@@
Finally, the APT specimens were prepared with the conventional focused ion beam scanning electron microscopy (FIB-SEM) in-situ lift-out protocol using a FIB-SEM dual beam instrument (FEI Versa 3D) with a single isotope Ga+ source (69 amu) 34, 35. An area of interest on the sample surface was identified and covered with a 2$\times$20 $\mu$m Pt protective strip deposited with the instrument’s gas injection system (GIS) using a (methyl-cyclopentadienyl)-trimethyl platinum precursor gas. A cantilever was prepared for lift out by milling trenches with the stage tilted 30° with respect to the ion beam. The cantilever was lifted out with an Omniprobe micromanipulator and placed in segments on a flat silicon or tungsten rod. The lift-out segments were polished with the focussed ion beam (FIB) by annular milling to obtain the needle-shaped sample shown in Figure 1b,c. We notice that the high charging effect due to the low electrical conductivity of silica reduces the image resolution.
@@ 2.2 Atom Probe Tomography
The LEAP 6000 XR operated in laser mode (wavelength 257.5 nm, $\approx$4.8 eV, deep-UV) with laser pulse energies (LPE) between 100 and 250 pJ focussed on a spot of 2 $\mu$m diameter, which corresponds to an energy density between 30 and 80 J/m2. The sample temperature was set to either 35 K or 50 K. The pulse rate was set using the instrument’s automatic pulse rate control so that ions could be detected up to 250 Da before the next pulse, regardless of the voltage. This meant that the pulse rate increased with increasing voltage due to shorter flight times.
@@ 2.3 Theoretical and computational methods
To produce defect-free amorphous silicon dioxide matrices, we assume ideal beta-cristobalite structures with a density of $0.0662$ atoms/Å3. The Monte Carlo (MC) approach developed by Wooten, Winer and Weaire 36 is then used to optimise the atomic configurations. In this method, the structure of the material is represented as a network of interconnected atoms, to which we first make a series of random bond changes, all of which are accepted. This results in a randomised structure, which was then annealed at constant volume by lowering the temperature at each step and performing a series of atomic displacements to optimise the bonding topology 37. The criterion for accepting (rejecting) such displacements is chosen according to the Boltzmann probability distribution, where the decrease (increase) in potential energy is evaluated using a force field parameterised for SiO2 38, similar to molecular dynamics simulations 39, 40, 41, 42. In this framework, we can amorphise the structures independently of their initial configurations and find minimum-energy configurations by annealing in a very efficient way. The resulting defect-free amorphous silica matrices are contained in a cubic box with periodic boundary conditions. In particular, we have generated three different sizes of simulation cells corresponding to a cubic box with an edge of $L=10.283$, $L=14.26$ and $L=21.39$ (see Figure S.3 of the SI for an image of one of the structures obtained).
The optical properties of these systems are evaluated in the framework of linear response theory by calculating the imaginary part of the complex dielectric tensor $\epsilon^{(2)}_{\alpha,\beta}$ (where $\alpha,\beta$ denotes the tensor components with respect to a set of Cartesian axes), whose spatial average is linearly related to the absorption coefficient (modulo some weighting factors). In our calculations, local field effects (LFE) are neglected and the expression for $\epsilon^{(2)}_{\alpha,\beta}$ is given by 24, 43
@@
where $e$ is the electron charge, $\Omega$ is the volume of the box, $N_{\mathbf{k}}$ is the number of points used to describe the electronic bands, $m$ is the mass of the electron, $\hat{\mathbf{M}}$ is the squared matrix element that connects the initial and final wave functions of the electrons via the radiation perturbative field, which belong to different bands labelled $n$ and $n^{\prime}$ but have the same crystal momentum $\mathbf{k}$, while $E_{n,\mathbf{k}}$ is the energy value of band $n$ at the point $\mathbf{k}$ of reciprocal space. Eq. 1 is obtained under the assumption that the transmitted momentum can be neglected (long wavelength limit), but can be extended to finite momentum transfer. The $\delta$ functions are slightly smeared to avoid unphysical excitations with infinite lifetimes. When comparing the spectra of the three boxes with different edges, we found that convergence is achieved with the box with $L=14.26$, which corresponds to 192 atoms (see Figure S.4 in the SI, where we compare the results for $L=10.28$, $L=14.26$ and $L=21.42$).
@@
The calculation of $\epsilon^{(2)}_{\alpha,\beta}$ was performed using DFT as implemented in the Quantum Espresso code suite 44. The ground state electron density was determined with a plane wave cutoff equal to $80$ Rydberg with the Perdew-Burke-Ernzerhof (PBE) functional. Using the same functional, we also investigated the effect of a finite constant electric field on the electronic properties of the pristine matrices and found negligible effects for all amorphous silica configurations. Once PBE self-consistency is achieved, we have improved the estimated electronic bandgap of pristine amorphous silica by using the HSE06 (Heyd–Scuseria–Ernzerhof) hybrid exchange–correlation functional HSE06 (Heyd–Scuseria–Ernzerhof) hybrid exchange–correlation functional 45 (see Figure S.5 in SI), which indeed yields a value of about 8 eV, which is much closer to the experimental values in the range of 8 to 9 eV 46, 47, 48.
@@
Recall that both sol-gel synthesis and specimen preparation for APT analysis usually result in defects in the pristine amorphous silica matrix, such as vacancies, dislocations and interstitials. In particular, the use of FIB for lift-out can lead to substitution or displacement of atomic constituents by the implantation of Ga+ ions, which can lose all their kinetic energy through elastic and inelastic interactions in the sample and be adsorbed. In the case of gallium, depending on the ion concentration, this can lead to significant changes in the structural and optical properties of the sample. To determine the effects of such defects on the optical properties of the sample, we performed the calculation of the imaginary part of the dielectric function in the long wavelength limit (see equation 1) for a number of typical defect configurations by including interstitial hydrogen, carbon and sodium atoms, oxygen vacancies 22, interstitial and substitutional gallium ions at different concentrations. We emphasise that only the positions of the atoms of the defective silica structures were optimised, while the periodic cells were constrained by density; therefore the cubic box has fixed lattice vectors. To optimise the atom positions, we use a BFGS quasi-Newton algorithm that enforces convergence to 10-5 Hartree/Bohr and 10-6 Hartree for the forces and energy, respectively. The optical properties for each defected configuration were evaluated using DFT with the HSE06 exchange-correlation functional 49, which ensures a more accurate evaluation of the resulting absorption properties.
3 Results